# Mixture of Experts Using Tensor Products

### v1

@@ 3.2 Tensor, Tensor Product, Entangled Tensor
The tensor $\mathcal{A}$ is a multidimensional array of elements (called components) of $\mathbb{R}$, each denoted by its integer coordinates in the array; e.g., for a two-dimensional array, the component at position $i,j\in\mathbb{N}$ is denoted $A_{i,j}$. The order of a tensor is how many indices it has (e.g., a vector $v$ is a first-order tensor, a matrix $M$ is a second-order tensor, etc.)
@@
The tensor product $V\otimes W$ of two vector spaces $\mathcal{V}$ and $\mathcal{W}$ is a vector space to which is associated a bilinear map $V\times W\rightarrow V\otimes W$ that maps a pair (v, w), $v\in V,w\in W$ to an element of $V\otimes W$ denoted $v\otimes w$. We can create tensor product spaces by more than one application of tensor product, $\mathcal{H}=\mathcal{U}\otimes\mathcal{V}\otimes\mathcal{W}$, with arbitrary bracketing, since tensor product is associative. Tensor product space of the form is said to have tensor order of $n$.
@@
The $n$-order tensor product space $\otimes^{n}_{j=1}\mathcal{H}_{j}$ consists of vectors of the form $v=\otimes^{n}_{j=1}v_{j}$, where $v_{j}\in\mathcal{H}_{j}$, are called simple tensor. Vectors need to be represented as the sum of multiple simple tensors called entangled tensors:
$\sum_{k=1}^{r}\bigotimes_{j=1}^{n}\mathcal{H}_{j}=\sum^{r}_{k=1}\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes...\otimes\mathcal{H}_{n}$ (2)
where tensor rank $r$ is is the smallest number of simple tensors that sum up to $v$. For example, $\frac{\psi_{0}\otimes\phi_{0}+\psi_{1}\otimes\phi_{1}}{\sqrt{2}}$ is a tensor of rank 2.
3.3 Tensorized Vector using Engangled Tensor
Any training parameters $v\in\mathbb{R}^{d}$ can be expressed as an entangled tensor of rank $r$ and order $n$ by: $v=\sum\limits_{k=1}^{r}\bigotimes_{j=1}^{n}v_{jk}$. Here, $v_{jk}\in\mathbb{R}^{q}$, yielding a resultant vector $v$ of dimension $p=q^{n}$. Its storage requirements are efficiently managed, consuming only $rnq=O(rq\log p/q)$. If $q^{n}>d$, the excess part of the generated vector will be cut off. The number of vector parameters can be reduced from $d$ to $rn\sqrt{d}$. For example, when $d=512$, $q=8$, $n=3$ and $r=2$, the number of parameters of a vector can be reduced from 512 to 48.
@@ 4 Methods: TensorPoly
Figure 2: TensorPoly-I and TensorPoly-II. We illustrate how to reparametrized the LoRA matrix $\mathbb{R}^{625\times 5}$ with 4 tensor $\mathcal{A}\in\mathbb{R}^{3\times 5\times 5}$. In this case, the tensor rank $R=3$, tensor order $N=4$. For TensorPoly-I, the routing function $Z$ is designed to select which rank of the entangled tensor is activated for a given task. Conversely, TensorPoly-II introduces a more granular control by selecting tensor rank and tensor order. The Blue color indicates a non-trainable component and reddish color indicates a trainable component
In this section, we introduced the Poly model. This section delves into two variant models, TensorPoly-I and TensorPoly-II. TensorPoly-I employs a routing mechanism that assigns distribution scores to tensor ranks while TensorPoly-II assigns distribution scores to the finer-grained tensor order.
@@ 4.1 TLoRA
TLoRA reparamizes the LoRA using the tensor product.
@@
$\begin{split}h&=W_{0}x+s\cdot A(B)^{\top}x\\ &=W_{0}x+s\cdot\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\mathcal{A}_{i,k}\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\mathcal{B}_{i,k}x\end{split}$ (TLoRA)
where $R$ is the rank, $N$ is the order of entangled tensor respectively.
4.2 Polytropon (Poly)
Poly/ MHR addresses the multi-task problem by softly sharing latent experts across tasks. Each Poly layer contains 1) an inventory of latent experts =$\mathcal{M}=\{\phi_{1},\ldots,\phi_{m}\}$ with $|\mathcal{M}|\ll|\mathcal{T}|$; 2) a routing function $r(\cdot)$ that chooses which subset of the experts to combine for each task.
Each module corresponds to a LoRA adapter, where $\phi_{i}$ are its associated parameters $A^{(i)},B^{(i)}\in\mathbb{R}^{d\times r}$. $r(\cdot)$ is implemented as a task–module routing matrix $Z\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{M}|}$. $z_{\tau}=Z_{\tau,:}\in\mathbb{R}^{|\mathcal{M}|}$ is a routing vector of task ${\mathcal{T}_{\tau}}$, with cell $Z_{\tau,j}$ being the probability logits of using module $\phi_{j}$ for task $\mathcal{T}_{\tau}$ in the current layer. Differently from mixture-of-experts (fedus2021switch), which perform token-level top-$k$ routing, $Z$ converges to a binary matrix, defining a soft partition over modules. This is achieved by using a Gumbel-sigmoid distribution (jang2016categorical) during training, with $\hat{Z}_{\tau,j}\sim\texttt{Gumbel}(Z_{\tau,j})$. At each forward pass, Poly can be defined as :
$A^{\tau}=\sum_{i}\alpha_{i}A^{(i)};\,B^{\tau}=\sum_{i}\alpha_{i}B^{(i)}$ (Poly)
4.3 TensorPoly-I
As illustrated in Figure 2, the TensorPoly model innovatively incorporates the Tensorized Low-Rank Adaptation (TLoRA) as its core module. This approach adopts finer-grained tensors for training parameters, thereby enabling a more nuanced parameter manipulation. The routing matrix $Z\in\mathbb{R}^{|\mathcal{T}|\times R}$ plays a pivotal role in this framework, determining which rank within the entangled tensor is to be activated for a given task. Upon activation of the selected rank, the model performs a linear combination of the selected tensor, weighted by a factor $\alpha$:
@@ 4.3 TensorPoly-I
$A^{\tau}=\sum\limits_{k=1}^{R}\alpha\bigotimes_{i=1}^{N}\mathcal{A}_{i,k};B^{\tau}=\sum\limits_{k=1}^{R}\alpha\bigotimes_{i=1}^{N}\mathcal{B}_{i,k}$ (TensorPoly-I)
@@
where $\alpha=\frac{\hat{z}_{\tau,i}}{\sum_{j}\hat{z}_{\tau,j}}$. $\mathcal{A}_{i,k}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil}$ is a third-order tensor.
4.4 TensorPoly-II
@@
In TensorPoly-II, we implement a fine-grained routing function that selects both the tensor rank and tensor order, enabling precise control over the model’s adaptation to specific tasks. This sophisticated routing mechanism facilitates the selection of finer-grained tensor elements, which are then aggregated through a tensor product operation followed by a linear combination. The routing itself is conceptualized as a third-order tensor $Z\in\mathbb{R}^{|\mathcal{T}|\times R\times N}$, which offers an unprecedented level of granularity in directing the model’s focus across different ranks and orders of the tensor space.
$A^{\tau}=\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\alpha\mathcal{A}_{i,k};B^{\tau}=\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\alpha\mathcal{B}_{i,k}$ (TensorPoly-II)
@@
Each unit of $Z$ is $\alpha=\frac{\hat{z}_{\tau,i,k}}{\sum_{j}\hat{z}_{\tau,j,k}}$, which will be routed to the specific order and rank in the entangled tensor.
@@ 5 Experiments
Model Natural Language Inference Sentence Completion Co-reference WSD ACC Param
@@
Baselines (w/o pretrain)
FullFT 79.8 87.5 46.6 41.3 40.0 81.0 46.4 93.8 65.4 56.5 57.7 63.3 3B
BitFit (with LN) 72.2 57.1 36.5 35.3 36.6 75.0 29.5 88.6 61.5 56.6 51.7 54.6 1.3M
LayerNorm 71.8 57.1 36.5 35.1 36.3 76.0 29.6 88.7 63.5 49.4 52.2 54.2 250K
Adapter 76.2 87.5 45.1 40.4 35.3 84.0 41.9 91.7 65.4 54.7 55.5 61.6 12.9M
Compacter 75.8 82.1 40.8 37.4 35.8 84.0 46.4 93.5 64.4 55.5 55.2 61.0 807K
Compacter++ 76.9 82.1 41.7 38.3 36.9 86.0 46.3 93.5 65.4 55.1 54.1 61.5 540K
Prompt(10) 52.7 66.1 34.2 33.5 33.5 67.0 29.9 84.2 54.8 51.9 51.6 50.9 41K
Prompt(100) 48.0 53.6 33.4 33.8 33.3 60.0 26.8 74.0 60.6 51.1 50.0 47.7 409K
Prefix tuning 68.6 84.0 43.3 37.5 36.5 71.0 42.1 90.2 56.7 52.0 54.2 57.8 576K
FishMask (0.2%) 76.9 83.9 43.7 39.7 37.2 82.0 44.1 94.2 63.5 54.5 52.5 61.1 6M
FishMask (0.02%) 75.5 76.8 39.9 38.1 36.2 84.0 38.2 93.6 61.5 53.9 53.5 59.2 600K
SAID 69.0 80.4 40.4 35.4 35.5 77.0 36.7 89.3 61.5 52.7 55.0 57.5 500K
SAID 66.1 83.9 41.3 38.5 35.8 76.0 38.3 89.7 55.8 50.9 55.3 57.4 20K
LoRA 78.3 85.7 45.1 41.0 39.5 88.0 47.1 93.6 60.6 56.8 55.2 62.8 9.1M
$\mathrm{(IA)}^{3}$ 78.0 87.5 48.6 40.8 40.8 87.0 49.4 94.7 68.3 59.8 56.0 64.6 540K
DoRA
w/ pretrain
LoRA 81.9 89.3 41.2 40.3 41.3 93.7 59.8 96.2 66.0 67.9 56.8 66.8 2.2M
$\mathrm{(IA)}^{3}$ 82.2 89.9 45.8 41.6 41.2 91.7 53.4 94.2 70.8 63.3 53.9 66.2 540K
DoRA
TLoRA 80.7 90.5 39.9 40.9 41.2 93.0 54.4 95.3 66.3 67.4 57.3 66.1 1.4M
Poly 84.7 89.3 46.0 42.8 42.7 93.0 63.3 96.6 68.9 70.1 59.9 68.8 17M
MHR 85.2 90.5 44.7 42.3 42.8 94.7 63.3 96.7 70.5 70.6 59.8 69.2 17M
TensorPoly-I 85.2 91.7 45.0 42.5 42.5 96.7 63.1 96.6 68.6 69.8 60.6 69.3 27.8M
TensorPoly-II 84.7 90.5 44.4 41.0 42.0 94.3 58.7 95.7 67.6 68.7 59.9 68.0 33.2M
Table 1: Results on the T0 few-shot benchmark. All the results in our implementation are the median score of 3 seeds [0,1024,42]. For all the baseline scores, we report the results from tfew. The bold is the best score. The underline is the second best in the same training setting.
Our experiments aim to answer the following questions: 1) How are different PEFT approaches performed? 2) Is routing important in multi-task scenarios? 3) Can non-linear merging methods in TensorPoly generalize well?
5.1 Backbone, Datasets and Evaluation
Backbone
To ensure our model retains high performance with a limited number of labeled examples after fine-tuning, it is crucial to select an appropriate pre-trained model as the backbone. To facilitate a fair comparison with baseline methodologies, we have chosen the T0 model, consistent with the approach described in the IA3 paper by tfew. T0 was created by finetuning the T5 model on a multi-task mixture of datasets (t5). Each dataset is associated with multiple prompt templates that are used to format the example to (input, target) pairs. Examples in the datasets to train the T0 were prompted by applying the prompt templates from the Public Pool of Prompts (P3) (bach2022promptsource). As a result, all the different types of natural language tasks can trained with sequence-to-sequence structure (t0_bench).

### v2

@@ 3.3 Tensor, Tensor Product, Entangled Tensor
The tensor $\mathcal{A}$ is a multi-dimensional array of elements (called components) of $\mathbb{R}$, each being denoted by its integer coordinates in the array; e.g., for a two-dimensional array, the component at position $i,j\in\mathbb{N}$ is denoted $A_{i,j}$. The order of a tensor is how many indices it has (e.g., a vector $v$ is a first-order tensor, a matrix $M$ is a second-order tensor, etc.).
@@
The tensor product $\mathcal{V}\otimes\mathcal{W}$ of two vector spaces $\mathcal{V}$ and $\mathcal{W}$ is a vector space to which is associated a bilinear map $\mathcal{V}\times\mathcal{W}\rightarrow\mathcal{V}\otimes\mathcal{V}$ that maps a pair of vectors $(v,w),v\in\mathcal{V},w\in\mathcal{W}$ to a vector in $\mathcal{V}\otimes\mathcal{W}$, denoted as $v\otimes w$. We can create tensor product spaces by more than one application of a tensor product, $\mathcal{H}=\mathcal{U}\otimes\mathcal{V}\otimes\mathcal{W}$, with arbitrary bracketing since the tensor product is associative. The tensor product space of $N$ vector spaces in such form is said to have a tensor order of $N$.
@@
The vectors in $N$-order tensor product space $\otimes^{N}_{j=1}\mathcal{H}_{j}$ are in the form $v=\otimes^{N}_{j=1}v_{j},v_{j}\in\mathcal{H}_{j}$, and referred to as simple tensors. Vectors represented as the sum of multiple simple tensors are called entangled tensors:
$\sum_{k=1}^{R}\bigotimes_{j=1}^{N}v_{jk}=\sum^{R}_{k=1}v_{1k}\otimes v_{2k}\otimes...\otimes v_{Nk}$ (2)
where tensor rank $R$ is the smallest number of simple tensors that sum up to $v$. For example, $\frac{v_{11}\otimes v_{21}+v_{12}\otimes v_{22}}{\sqrt{2}}$ is a tensor of rank 2.
3.4 Tensorized Vector using Entangled Tensor
Any training parameter $v\in\mathbb{R}^{d}$ can be expressed as an entangled tensor of rank $R$ and order $N$ by: $v=\sum\limits_{k=1}^{R}\bigotimes_{j=1}^{N}v_{jk}$. Here, $v_{jk}\in\mathbb{R}^{Q}$, yielding a resultant vector $v$ of dimension $p=q^{N}$. If there exists a value of $N$ such that $d=p$, the storage requirements are efficiently managed, consuming only $RNq=O(Rq\log p/q)$ parameters. More generally, we take the smallest possible value $q$ satisfying $q^{N}>d$, i.e. $q=\lceil\sqrt[N]{d}\rceil$, and cut off the excess part of the generated vector to produce $v\in\mathbb{R}^{d}$.
For example, we would like to compress a 512-dim vector ($d=512$) with a tensor of order $N=3$ and rank $R=2$. Then we can set the dimension of each small vectors as $q=\sqrt[N]{d}=8$. In this way, the vector can be represented with a total number of $RNq=48$ parameters, leading to a significant parameter reduction.
@@ 4 Methods: TensorPoly
We propose two variant TensorPoly models: TensorPoly-I and TensorPoly-II based on the TLoRA module (§$\S$ 4.1). TensorPoly-I employs a routing mechanism that assigns distribution scores to tensor ranks (§$\S$ 4.2) while TensorPoly-II assigns distribution scores to the finer-grained tensor order.
@@ 4.1 TLoRA
[…]
@@
$\begin{split}h&=W_{0}x+s\cdot A(B)^{\top}x\\ &=W_{0}x+s\cdot\left(\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\mathcal{A}_{i,k}\right)\left(\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\mathcal{B}_{i,k}\right)^{T}x\end{split}$ (TLoRA)
where $R$ is the tensor rank, $N$ is the order of entangled tensor respectively. $\mathcal{A},\mathcal{B}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil\times R)}$ are fourth-order tensors, which refer to the training parameters of the modules. $\mathcal{A}_{i,k},\mathcal{B}_{i,k}\in\mathbb{R}^{r\times\lceil\sqrt[N]{d}\rceil)}$ are indices of fourth-order tensors, referring to specific rank and order in an entangled tensor, as depicted in Fig 3.
4.2 TensorPoly: Mixture of Latent Experts using Tensor Products
Figure 3: TensorPoly-I and TensorPoly-II. We illustrate how to reparameterize the LoRA matrix $\mathbb{R}^{625\times 5}$ with 4 tensor $\mathcal{A}\in\mathbb{R}^{3\times 5\times 5}$. In this case, the tensor rank $R=3$, tensor order $N=4$. For TensorPoly-I, the routing function $\mathbf{Z}$ is designed to select which rank of the entangled tensor is activated for a given task. Conversely, TensorPoly-II introduces a more granular control by selecting tensor rank and tensor order.
Latent-expert approaches have been proven effective in the few-shot multi-task transfer learning (Ponti et al., 2023; Caccia et al., 2023). Poly merges the latent experts by averaging the weight. MHR partitioned the LoRA into several heads and use a piecewise linear aggregation (i.e., linear within disjoint intervals). Instead of using a linear combination, we propose TensorPoly models that incorporate TLoRA as the core module. Specifically, we propose a TensorPoly-I and TensorPoly-II according to the routing granularity. In each layer, TensorPoly-I use a routing matrix $Z\in\mathbb{R}^{|\mathcal{T}|\times R}$ to determine which rank within the entangled tensor can be activated for a given task. In this case, each rank $\mathcal{A}_{k}$ corresponds to an “expert” in the forward pass. Upon activation of the selected rank, the model computes a linear combination of the selected rank, weighted by a factor $\alpha$:
@@ 4.2 TensorPoly: Mixture of Latent Experts using Tensor Products
$\begin{split}\mathcal{A}_{k}&=\bigotimes_{i=1}^{N}\mathcal{A}_{i,k}=\mathcal{A}_{1,k}\otimes\mathcal{A}_{2,k}\otimes...\otimes\mathcal{A}_{N,k}\\ A^{\tau}&=\sum\limits_{k=1}^{R}\alpha_{k}\mathcal{A}_{k}=\underbrace{\alpha_{1}\mathcal{A}_{1}+...+\alpha_{R}\mathcal{A}_{R}}_{\textbf{merge the rank $R$}}\end{split}$ (TensorPoly-I)
@@
The same for $B^{\tau}=\sum\limits_{k=1}^{R}\alpha_{k}\mathcal{B}_{k}$, where $\alpha_{k}=\frac{\hat{z}_{\tau,k}}{\sum_{j}\hat{z}_{\tau,j}}$. The expert $\mathcal{A}_{k}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil}$ is a third-order tensor.
@@
Contrast to the TensorPoly-I with a routing matrix, the routing for TensorPoly-II is conceptualized as a third-order tensor $\mathbf{Z}\in\mathbb{R}^{|\mathcal{T}|\times N\times R}$, which offers a finer-grained level in directing the model’s focus across different ranks and orders of the tensor space. This sophisticated routing mechanism facilitates the selection of finer-grained tensor elements, which are then aggregated through a tensor product operation.
$\begin{split}A_{i}^{\tau}&=\sum\limits_{k=1}^{R}\alpha_{i,k}\mathcal{A}_{i,k}=\underbrace{\alpha_{i,1}\mathcal{A}_{i,1}+...+\alpha_{i,R}\mathcal{A}_{i,R}}_{\textbf{merge the rank $R$}}\\ A^{\tau}&=\bigotimes_{i=1}^{N}\mathcal{A}_{i}^{\tau}=\underbrace{\mathcal{A}_{1}^{\tau}\otimes\mathcal{A}_{2}^{\tau}\otimes...\otimes\mathcal{A}_{N}^{\tau}}_{\textbf{merge the order $N$}}\end{split}$ (TensorPoly-II)
@@
Each unit of $Z$ is $\alpha_{i,k}=\frac{\hat{z}_{\tau,i,k}}{\sum_{j}\hat{z}_{\tau,j,k}}$, which will be routed to the specific order and rank in the entangled tensor. In this case, an “expert” $\mathcal{A}_{i,k}\in\mathbb{R}^{r\times\lceil\sqrt[N]{d}\rceil}$ corresponds to each order in the entangled tensor.
@@ 5 Experiments
Model Natural Language Inference Sentence Completion Co-reference WSD ACC
@@
Baselines (w/o pre-train)
FullFT 79.8 87.5 46.6 41.3 40.0 81.0 46.4 93.8 65.4 56.5 57.7 63.3
BitFit (with LN) 72.2 57.1 36.5 35.3 36.6 75.0 29.5 88.6 61.5 56.6 51.7 54.6
LayerNorm 71.8 57.1 36.5 35.1 36.3 76.0 29.6 88.7 63.5 49.4 52.2 54.2
Adapter 76.2 87.5 45.1 40.4 35.3 84.0 41.9 91.7 65.4 54.7 55.5 61.6
Compacter 75.8 82.1 40.8 37.4 35.8 84.0 46.4 93.5 64.4 55.5 55.2 61.0
Compacter++ 76.9 82.1 41.7 38.3 36.9 86.0 46.3 93.5 65.4 55.1 54.1 61.5
Prompt(10) 52.7 66.1 34.2 33.5 33.5 67.0 29.9 84.2 54.8 51.9 51.6 50.9
Prompt(100) 48.0 53.6 33.4 33.8 33.3 60.0 26.8 74.0 60.6 51.1 50.0 47.7
Prefix tuning 68.6 84.0 43.3 37.5 36.5 71.0 42.1 90.2 56.7 52.0 54.2 57.8
FishMask (0.2%) 76.9 83.9 43.7 39.7 37.2 82.0 44.1 94.2 63.5 54.5 52.5 61.1
FishMask (0.02%) 75.5 76.8 39.9 38.1 36.2 84.0 38.2 93.6 61.5 53.9 53.5 59.0
SAID 69.0 80.4 40.4 35.4 35.5 77.0 36.7 89.3 61.5 52.7 55.0 57.5
LoRA 78.3 85.7 45.1 41.0 39.5 88.0 47.1 93.6 60.6 56.8 55.2 62.8
$\mathrm{(IA)}^{3}$ 78.0 87.5 48.6 40.8 40.8 87.0 49.4 94.7 68.3 59.8 56.0 64.6
w/ pre-train
LoRA 81.9 89.3 41.2 40.3 41.3 93.7 59.8 96.2 66.0 67.9 56.8 66.8
$\mathrm{(IA)}^{3}$ 82.2 89.9 45.8 41.6 41.2 91.7 53.4 94.2 70.8 63.3 53.9 66.2
TLoRA 80.7 90.5 39.9 40.9 41.2 93.0 54.4 95.3 66.3 67.4 57.3 66.1
Poly 84.7 89.3 46.0 42.8 42.7 93.0 63.3 96.6 68.9 70.1 59.9 68.8
MHR 85.2 90.5 44.7 42.3 42.8 94.7 63.3 96.7 70.5 70.6 59.8 69.2
TensorPoly-I 85.2 91.7 45.0 42.5 42.5 96.7 63.1 96.6 68.6 69.8 60.6 69.3
TensorPoly-II 84.7 90.5 44.4 41.0 42.0 94.3 58.7 95.7 67.6 68.7 59.9 68.0
Table 1: Results on the T0 few-shot benchmark. All the results in our implementation are the median score of 3 random seeds [0, 1024, 42]. For all the baseline scores, we quote the results from Liu et al. (2022). The value in bold is the best score.
To evaluate the effectiveness of our approaches, we perform experiments on multi-task transfer learning datasets T0 benchmark (Sanh et al., 2021), which is widely used in few-shot generalization approaches. In addition, a diverse array of tasks in T0 benchmark can help us test the generalization ability across different tasks. We conduct a comparative analysis between routing approaches (Poly, MHR, TensorPoly-I, TensorPoly-II) and their corresponding “single-expert” (without routing function) version (LoRA,TLoRA), as detailed in §$\S$5.2. We also investigate how parameter efficiency and effectiveness are influenced by hyperparameters rank $R$ and order $N$ in an entangled tensor (§5.3.1).
5.1 Datasets and Evaluation